# QOCO: A Quadratic Objective Conic Optimizer with Custom Solver Generation

### v2

@@ Appendix D Detailed Numeric Results
LASER 9216 10 10 - 21 24 0.00258 0.00414 - 0.02421 0.03150
QPCBOEI1 4743 21 15 - 21 33 0.00613 0.00632 - 0.01534 0.03697
HUESTIS 39968 7 8 - 14 - 0.01261 0.02198 - 0.03369 -
QBANDM 3007 21 20 32 29 22 0.00345 0.00502 0.01198 0.01520 0.02451
QSC205 775 15 18 17 30 15 0.00111 0.00192 0.00183 0.00950 0.00823
LISWET11 40002 - - - 40 31 - - - 0.15115 0.26106
LISWET8 40002 - - - 417 70 - - - 1.83018 0.90131
AUG2DC 60200 0 0 - 2 15 0.01392 0.02377 - 0.02266 0.46025
QSEBA 6548 30 29 - 27 30 0.00717 0.01452 - 0.02000 0.05188
AUG2D 59800 0 0 - 2 11 0.01088 0.02133 - 0.01970 0.30137
QGROW15 7365 18 20 17 20 28 0.00616 0.01114 0.01225 0.02356 0.06942
CONT-200 319196 13 10 30 12 28 3.41491 3.91965 8.99180 0.75313 2.27627
LISWET1 40002 - - - 27 52 - - - 0.09510 0.55497
ZECEVIC2 9 6 7 8 8 7 0.00002 0.00005 0.00004 0.00275 0.00102
QFFFFF80 8997 95 25 45 48 22 0.08216 0.03295 0.49989 0.06666 0.10850
AUG2DCQP 80400 11 12 - 17 - 0.08341 0.14402 - 0.09956 -
QPILOTNO 16258 39 31 43 56 27 0.09448 0.11113 0.21933 0.12265 0.15589
CVXQP2_M 6733 13 9 30 15 17 0.03255 0.02544 11.03589 0.06452 1.07815
POWELL20 30000 38 - - 60 - 0.05451 - - 0.21773 -
STADAT2 15997 17 44 - 18 21 0.00939 0.04388 - 0.02200 0.04019
QPCBOEI2 1623 40 19 - 26 48 0.00298 0.00274 - 0.00741 0.01278
PRIMAL3 22292 11 8 18 17 8 0.01240 0.01680 0.03449 0.01265 0.04159
QSHIP08S 21178 14 15 26 24 18 0.02043 0.02884 1.58040 0.04072 0.17699
DUALC2 1645 18 10 13 11 16 0.00054 0.00061 0.00091 0.00860 0.00116
LISWET4 40002 10 22 - 13 19 0.01575 0.05955 - 0.05271 0.13555
QPCSTAIR 4872 25 22 - 23 39 0.00897 0.01026 - 0.02300 0.05212
PRIMALC2 2076 24 15 30 11 16 0.00130 0.00142 0.00353 0.00385 0.00828
LISWET6 40002 9 17 - 13 20 0.01906 0.05229 - 0.05445 0.14808
QSHIP04L 8506 14 15 26 21 28 0.00769 0.00918 0.02240 0.02517 0.06894
QE226 3824 18 24 22 25 19 0.00466 0.00709 0.02897 0.01216 0.01492
PRIMAL4 17520 9 9 16 16 11 0.01266 0.01348 0.02799 0.00947 0.04455
QSCSD8 13844 11 10 15 15 12 0.00735 0.01020 0.05332 0.02534 0.06593
QSCORPIO 1824 15 11 13 24 15 0.00128 0.00169 0.00258 0.00776 0.03019
CONT-201 290393 12 10 - 14 - 4.17722 5.71965 - 0.82300 -
GOULDQP2 3142 9 12 20 13 10 0.00188 0.00313 0.58865 0.00962 0.05237
LISWET10 40002 - - - 125 20 - - - 0.44295 0.14301
HS268 40 5 10 15 12 15 0.00003 0.00007 0.00009 0.00356 0.00057
QBORE3D 1834 18 28 24 26 20 0.00202 0.00442 0.00487 0.00950 0.01198
QSHIP12L 83825 18 15 34 23 39 0.21193 0.19337 79.39230 0.10314 4.08685
QSHIP04S 5866 14 14 24 19 21 0.00373 0.00592 0.01427 0.01438 0.04390
QSHARE1B 1415 28 28 46 30 31 0.00200 0.00343 0.00512 0.00929 0.02247
QGFRDXPN 3889 24 20 - 43 31 0.00589 0.00682 - 0.03540 0.08180
CONT-050 19796 10 9 24 10 18 0.03964 0.04369 0.10372 0.02808 0.08417
MOSARQP1 8467 9 10 24 15 11 0.00520 0.00843 0.03596 0.01658 0.02250
BOYD2 517049 68 - - 78 - 2.99925 - - 6.10176 -
QSCAGR25 2182 17 19 34 24 - 0.00160 0.00345 0.00844 0.01023 -
HUES-MOD 39968 13 12 - 13 23 0.02296 0.02834 - 0.02918 0.36701
YAO 8007 - - - 25 36 - - - 0.01411 0.05955
CONT-101 72693 10 10 - 13 25 0.30755 0.44878 - 0.09657 0.47646
HS53 24 5 6 9 8 7 0.00003 0.00004 0.00005 0.00273 0.00030
CVXQP1_M 7482 13 10 - 16 21 0.04223 0.03895 - 0.09931 0.43868
STCQP2 48135 7 8 25 12 14 0.08184 0.10365 0.53351 0.21152 0.16697
QGROW22 10930 22 25 20 25 31 0.00986 0.01891 0.02126 0.03386 0.10940
HS35 11 5 7 10 9 9 0.00002 0.00005 0.00004 0.00366 0.00077
STCQP1 48135 7 7 28 - 24 0.03555 0.04303 0.32154 - 0.24021
STADAT1 15997 20 15 - 185 - 0.01086 0.01243 - 0.17253 -
PRIMAL1 6140 12 9 17 18 8 0.00325 0.00402 0.00986 0.00567 0.00878
DUAL3 6441 9 11 13 11 7 0.00313 0.00378 0.00485 0.01875 0.00723
HS21 8 9 8 8 8 10 0.00002 0.00003 0.00003 0.00255 0.00027
QBEACONF 3664 12 - 29 18 22 0.00182 - 0.00717 0.00796 0.01373
QISRAEL 3109 26 25 31 23 23 0.00366 0.00576 0.01015 0.01732 0.00747
LISWET2 40002 9 17 - 15 20 0.01800 0.04961 - 0.05521 0.14195
DTOC3 49994 4 6 - 8 26 0.01081 0.02869 - 0.02627 0.38133
GOULDQP3 3840 6 7 25 9 17 0.00109 0.00237 3.86470 0.01007 0.17719
PRIMAL2 8691 11 8 16 17 8 0.00431 0.00527 0.01403 0.00586 0.01458
TAME 7 4 5 5 6 4 0.00002 0.00004 0.00003 0.00368 0.00031
PRIMALC5 2860 20 13 14 11 14 0.00141 0.00149 0.00241 0.00348 0.00938
AUG3DC 10419 0 0 14 2 10 0.00227 0.00270 0.03786 0.00950 0.04388
AUG3DCQP 14292 9 10 29 16 22 0.01343 0.01937 0.08650 0.02952 0.16977
QSCFXM2 7228 32 30 36 56 - 0.01022 0.01737 0.04520 0.05266 -
VALUES 4428 12 13 - - - 0.00149 0.00242 - - -
QAFIRO 121 15 13 14 15 13 0.00017 0.00023 0.00020 0.00347 0.00100
QETAMACR 7761 26 19 25 28 27 0.05831 0.04680 0.91543 0.04582 0.10767
QSCSD6 7070 14 13 16 17 13 0.00407 0.00613 0.02588 0.01711 0.05499
QSTAIR 5423 23 33 32 41 40 0.01108 0.01966 0.03227 0.03192 0.05891
PRIMALC1 2514 25 16 27 10 20 0.00237 0.00153 0.00347 0.00359 0.01178
QSHELL 40488 28 36 - 33 27 0.10774 0.12657 - 0.07681 0.18787
CVXQP2_L 67483 22 9 - 16 31 15.34297 9.37783 - 2.75816 750.19697
QSCFXM1 3779 26 25 35 49 31 0.00448 0.00762 0.02215 0.02710 0.08131
DUAL1 3813 10 12 14 12 12 0.00182 0.00219 0.00283 0.01026 0.00219
UBH1 66033 7 17 - 16 8 0.08430 0.08050 - 0.05857 0.15917
DUALC5 2276 10 10 18 14 12 0.00043 0.00084 0.00126 0.01239 0.00120
LISWET9 40002 - - - - 62 - - - - 0.75226
Q25FV47 71470 27 25 34 37 53 0.18642 0.18053 2.34299 0.16627 0.52865
PRIMALC8 5182 20 12 14 10 22 0.00247 0.00322 0.00401 0.00440 0.02708
QPCBLEND 657 16 16 17 20 17 0.00060 0.00095 0.00097 0.00530 0.00192
LOTSCHD 72 7 8 18 11 15 0.00004 0.00008 0.00011 0.00273 0.00067
QSCAGR7 585 15 15 29 23 32 0.00044 0.00084 0.00174 0.00461 0.00756
QSCRS8 4472 23 31 33 35 30 0.00722 0.01372 0.01325 0.02854 0.06120
CVXQP1_S 734 8 8 12 13 12 0.00054 0.00085 0.01330 0.00830 0.00374
QADLITTL 567 11 13 17 16 14 0.00041 0.00076 0.00140 0.00485 0.00326
CVXQP2_S 660 8 9 12 14 11 0.00045 0.00073 0.01244 0.00789 0.00310
QRECIPE 988 18 16 19 20 15 0.00110 0.00153 0.00195 0.00703 0.00469
QSCSD1 3893 11 10 12 16 12 0.00267 0.00363 0.00896 0.01347 0.01324
GENHS28 43 0 0 9 1 5 0.00002 0.00004 0.00008 0.00576 0.00039
QSCTAP1 2325 15 10 10 17 13 0.00176 0.00185 0.00272 0.00821 0.02044
QBRANDY 2462 19 18 28 40 28 0.00330 0.00420 0.00662 0.01947 0.02632
HS52 14 0 0 7 2 8 0.00002 0.00003 0.00005 0.00333 0.00039
HS51 14 0 0 7 2 7 0.00001 0.00002 0.00004 0.00242 0.00025
CVXQP3_S 808 9 10 13 14 14 0.00050 0.00089 0.01110 0.00681 0.00338
CVXQP3_L 82481 - 10 - 14 42 - 21.86824 - 6.19038 187.95385
STADAT3 31997 17 54 28 27 22 0.02096 0.10802 0.08724 0.06329 0.07766
S268 40 5 10 15 12 15 0.00002 0.00006 0.00007 0.00266 0.00044
QSTANDAT 5030 24 20 11 47 22 0.00536 0.00792 0.04844 0.03088 0.05748
EXDATA 1137750 16 21 19 14 13 4.82029 6.01972 36.06845 1.83909 0.83910
QGROW7 3550 19 21 16 20 27 0.00310 0.00556 0.00557 0.01416 0.03009
DUAL4 3024 8 11 13 10 13 0.00134 0.00224 0.00236 0.00856 0.00247
QFORPLAN 5591 28 21 - 56 - 0.00617 0.00799 - 0.02855 -
QSCTAP3 12401 16 11 11 19 12 0.01999 0.01361 0.07613 0.03330 0.08330
QSHARE2B 828 19 15 24 21 20 0.00065 0.00092 0.00141 0.00396 0.00258
HS35MOD 12 9 11 11 15 13 0.00002 0.00003 0.00005 0.00227 0.00032
DUAL2 4796 8 10 13 11 7 0.00168 0.00263 0.00429 0.00878 0.00594
MOSARQP2 4775 8 9 27 16 13 0.00445 0.00614 0.02766 0.01492 0.01722
QPTEST 10 5 7 9 10 10 0.00001 0.00003 0.00003 0.00225 0.00021
AUG3D 9219 0 0 13 2 6 0.00191 0.00252 0.03937 0.00711 0.04820
QSIERRA 11557 22 26 38 24 21 0.01371 0.02585 0.05065 0.03655 0.12325
LISWET5 40002 8 9 - 13 39 0.01421 0.03237 - 0.05394 0.40437
LISWET3 40002 9 16 - 13 20 0.01809 0.05045 - 0.05186 0.14603
LISWET12 40002 - - - - 60 - - - - 0.72638
CVXQP3_M 8231 23 11 - 18 21 0.08985 0.05016 - 0.14954 0.46277
HS118 108 11 11 9 13 10 0.00007 0.00012 0.00011 0.00229 0.00048
QSHIP12S 28344 18 15 31 24 26 0.02525 0.03026 7.90644 0.04829 0.79477
DUALC1 1998 23 11 29 15 14 0.00076 0.00074 0.00162 0.01125 0.00095
CVXQP1_L 74982 - 10 - 15 37 - 19.33309 - 4.44083 157.56906
LISWET7 40002 11 - - 58 44 0.01843 - - 0.26488 0.49234
CONT-100 79596 11 9 22 11 19 0.32731 0.38091 0.80000 0.08902 0.27225
QSCFXM3 10369 31 32 34 58 31 0.02114 0.02997 0.07188 0.08477 0.09320
DPKLO1 1652 0 0 13 2 4 0.00018 0.00027 0.00204 0.00283 0.00136
KSIP 18871 14 14 21 17 11 0.00666 0.00982 0.01018 0.01295 0.00380
DUALC8 4076 14 9 27 14 22 0.00111 0.00136 0.00493 0.02466 0.00201
QSHIP08L 52050 14 15 28 22 20 0.11742 0.14559 13.08404 0.08743 0.45644
AUG3DQP 13092 11 11 31 17 13 0.01592 0.02040 0.08446 0.03143 0.07889
AUG2DQP 80000 11 13 - 17 - 0.07453 0.14683 - 0.10080 -
HS76 20 6 6 10 10 7 0.00002 0.00004 0.00005 0.00300 0.00038
QSCTAP2 9371 14 10 10 18 11 0.00958 0.00923 0.03774 0.02440 0.07304
BOYD1 745507 37 - - 30 - 0.90575 - - 1.08528 -
QCAPRI 3147 40 31 - 33 39 0.00557 0.00889 - 0.02008 0.06086
CONT-300 653093 12 10 - 14 - 15.35224 15.73776 - 2.06562 -

### v3

@@ Appendix D Detailed numeric results
LASER 9216 10 10 - 21 24 0.00295 0.00418 - 0.02075 0.03384
QPCBOEI1 4743 21 15 - 22 44 0.00557 0.00623 - 0.00675 0.04344
HUESTIS 39968 7 8 - 14 - 0.01247 0.02049 - 0.01075 -
QBANDM 3007 21 20 30 21 17 0.00283 0.00447 0.00984 0.00504 0.00617
QSC205 775 15 18 17 17 22 0.00085 0.00158 0.00146 0.00349 0.00221
LISWET11 40002 - - - 40 31 - - - 0.13571 0.27070
LISWET8 40002 - - - 417 57 - - - 1.72564 0.69070
AUG2DC 60200 0 0 - 2 12 0.01347 0.01958 - 0.02400 0.23210
QSEBA 6548 30 29 - 20 34 0.00622 0.01452 - 0.00404 0.01449
AUG2D 59800 0 0 - 2 14 0.00958 0.01984 - 0.02289 0.31375
QGROW15 7365 18 20 17 17 22 0.00529 0.00997 0.00961 0.00758 0.01993
CONT-200 319196 13 10 30 12 24 3.08655 4.03513 8.48776 0.39624 1.77756
LISWET1 40002 - - - 27 45 - - - 0.09196 0.50075
ZECEVIC2 9 6 7 8 8 7 0.00003 0.00005 0.00004 0.00255 0.00106
QFFFFF80 8997 93 25 45 26 26 0.06986 0.02945 0.42187 0.01826 0.09688
AUG2DCQP 80400 11 12 - 17 - 0.08231 0.13210 - 0.05561 -
QPILOTNO 16258 38 31 39 26 29 0.07808 0.09577 0.32104 0.03651 0.11626
CVXQP2_M 6733 13 9 30 12 16 0.02790 0.02295 10.49264 0.02690 0.70517
POWELL20 30000 38 - - 60 - 0.05336 - - 0.20098 -
STADAT2 15997 17 44 - 18 21 0.00891 0.04415 - 0.01993 0.04982
QPCBOEI2 1623 40 19 - 27 35 0.00236 0.00273 - 0.00413 0.00854
PRIMAL3 22292 11 8 18 18 8 0.01258 0.01602 0.03150 0.01333 0.01050
QSHIP08S 21178 14 15 27 20 17 0.01986 0.02778 1.55641 0.02175 0.15395
DUALC2 1645 18 10 13 9 15 0.00053 0.00056 0.00085 0.00207 0.00080
LISWET4 40002 10 22 - 13 20 0.01496 0.05909 - 0.05490 0.15414
QPCSTAIR 4872 25 22 - 21 34 0.00806 0.01108 - 0.01309 0.02036
PRIMALC2 2076 24 15 30 10 12 0.00120 0.00119 0.00306 0.00213 0.00196
LISWET6 40002 9 17 - 13 21 0.01726 0.05086 - 0.04968 0.17387
QSHIP04L 8506 14 15 25 17 22 0.00711 0.00944 0.01949 0.00936 0.01942
QE226 3824 18 24 22 19 18 0.00422 0.00677 0.02736 0.00802 0.01922
PRIMAL4 17520 9 9 16 16 9 0.01190 0.01330 0.02602 0.00949 0.01691
QSCSD8 13844 11 10 15 14 10 0.00644 0.00955 0.05031 0.01281 0.03718
QSCORPIO 1824 15 11 13 15 14 0.00122 0.00165 0.00236 0.00320 0.00257
CONT-201 290393 12 10 - 12 - 3.68995 5.09101 - 0.37022 -
GOULDQP2 3142 9 12 20 11 7 0.00177 0.00315 0.56048 0.00697 0.04653
LISWET10 40002 - - - 125 20 - - - 0.40801 0.15746
HS268 40 5 10 15 12 15 0.00003 0.00007 0.00009 0.00301 0.00053
QBORE3D 1834 18 28 24 11 15 0.00198 0.00440 0.00724 0.00253 0.00180
QSHIP12L 83825 18 15 32 18 29 0.20070 0.19269 72.93811 0.05007 2.43260
QSHIP04S 5866 14 14 25 16 20 0.00358 0.00600 0.01307 0.00542 0.01263
QSHARE1B 1415 28 28 43 21 26 0.00190 0.00319 0.00469 0.00467 0.02870
QGFRDXPN 3889 24 20 - 19 - 0.00537 0.00846 - 0.00708 -
CONT-050 19796 10 9 24 10 18 0.03774 0.04240 0.10226 0.01759 0.08744
MOSARQP1 8467 9 10 24 16 12 0.00509 0.00884 0.03554 0.01028 0.01954
BOYD2 517049 68 - - 90 - 1.79598 - - 0.87993 -
QSCAGR25 2182 17 19 34 18 31 0.00173 0.00351 0.00796 0.00450 0.01722
HUES-MOD 39968 13 12 - 13 18 0.02211 0.02751 - 0.01053 0.07958
YAO 8007 - - - 28 34 - - - 0.01528 0.05277
CONT-101 72693 10 10 - 11 80 0.30443 0.43120 - 0.08288 1.72033
HS53 24 5 6 9 8 6 0.00003 0.00005 0.00007 0.00261 0.00047
CVXQP1_M 7482 13 10 - 14 22 0.04193 0.03792 - 0.05341 0.45150
STCQP2 48135 7 8 25 10 11 0.07536 0.10185 0.51387 0.03917 0.04279
QGROW22 10930 22 25 20 19 24 0.00963 0.01915 0.01957 0.01047 0.04234
HS35 11 5 7 10 9 9 0.00002 0.00005 0.00003 0.00266 0.00999
STCQP1 48135 7 7 28 9 11 0.03430 0.04184 0.31809 0.00664 0.02806
STADAT1 15997 20 15 - - - 0.01204 0.01224 - - -
PRIMAL1 6140 12 9 17 18 11 0.00330 0.00377 0.00911 0.00539 0.00803
DUAL3 6441 9 11 13 15 7 0.00287 0.00433 0.00484 0.00864 0.00370
HS21 8 9 8 8 0 11 0.00002 0.00003 0.00003 0.00205 0.00034
QBEACONF 3664 12 - 29 11 14 0.00182 - 0.00648 0.00268 0.00234
QISRAEL 3109 26 25 30 21 22 0.00340 0.00542 0.00873 0.01089 0.01127
LISWET2 40002 9 17 - 15 20 0.01736 0.04994 - 0.05463 0.15266
DTOC3 49994 4 6 - 2 14 0.01034 0.02675 - 0.01440 0.13182
GOULDQP3 3840 6 7 25 8 16 0.00103 0.00196 3.92135 0.00670 0.17622
PRIMAL2 8691 11 8 16 17 7 0.00437 0.00506 0.01278 0.00645 0.00473
TAME 7 4 5 5 6 4 0.00002 0.00004 0.00003 0.00259 0.00034
PRIMALC5 2860 20 13 14 11 11 0.00141 0.00145 0.00216 0.00210 0.00236
AUG3DC 10419 0 0 14 2 7 0.00201 0.00302 0.03647 0.00928 0.01731
AUG3DCQP 14292 9 10 29 16 10 0.01278 0.01802 0.08718 0.01828 0.02658
QSCFXM2 7228 32 30 37 31 37 0.01037 0.01937 0.04069 0.01564 0.06499
VALUES 4428 11 13 - - - 0.00165 0.00226 - - -
QAFIRO 121 15 13 15 12 14 0.00015 0.00021 0.00018 0.00265 0.00661
QETAMACR 7761 26 19 26 21 21 0.05323 0.04463 0.94189 0.01268 0.06093
QSCSD6 $70$ 14 13 16 15 13 0.00372 0.00569 0.02306 0.00688 0.04069
QSTAIR 5423 23 33 32 20 38 0.01015 0.01912 0.02920 0.00891 0.03350
PRIMALC1 2514 25 16 27 11 14 0.00216 0.00148 0.00321 0.00279 0.00321
QSHELL 40488 28 36 - 38 26 0.09449 0.12115 - 0.07993 0.18460
CVXQP2_L 67483 22 9 - 15 23 14.21961 8.60789 - 0.47604 390.09870
QSCFXM1 3779 26 25 36 23 35 0.00385 0.00677 0.02216 0.00790 0.03538
DUAL1 3813 10 12 14 15 12 0.00173 0.00259 0.00275 0.00672 0.00236
UBH1 66033 7 17 - 6 11 0.02002 0.07995 - 0.01160 0.05267
DUALC5 2276 10 10 18 11 12 0.00047 0.00071 0.00116 0.00244 0.00067
LISWET9 40002 - - - - 53 - - - - 0.63997
Q25FV47 71470 27 25 33 27 41 0.16469 0.17414 4.81007 0.08488 0.33967
PRIMALC8 5182 20 12 14 9 13 0.00247 0.00326 0.00375 0.00220 0.00417
QPCBLEND 657 16 16 17 19 20 0.00055 0.00088 0.00115 0.00355 0.00547
LOTSCHD 72 7 8 18 10 13 0.00004 0.00006 0.00012 0.00201 0.00053
QSCAGR7 585 15 15 29 18 49 0.00042 0.00081 0.00128 0.00274 0.00534
QSCRS8 4472 23 31 33 26 24 0.00487 0.01325 0.01523 0.00647 0.01403
CVXQP1_S 734 8 8 12 11 12 0.00050 0.00076 0.01201 0.00416 0.00402
QADLITTL 567 11 13 17 12 16 0.00037 0.00070 0.00128 0.00330 0.00193
CVXQP2_S 660 8 9 12 12 10 0.00035 0.00058 0.01285 0.00294 0.00394
QRECIPE 988 18 16 19 19 18 0.00111 0.00149 0.00185 0.00347 0.00222
QSCSD1 3893 11 10 12 13 9 0.00251 0.00348 0.00695 0.00614 0.00966
GENHS28 43 0 0 9 1 5 0.00001 0.00002 0.00006 0.00262 0.01276
QSCTAP1 2325 15 10 10 0 12 0.00173 0.00191 0.00247 0.00246 0.01201
QBRANDY 2462 19 18 27 17 14 0.00299 0.00430 0.00576 0.00514 0.00510
HS52 14 0 0 7 2 6 0.00001 0.00003 0.00004 0.00266 0.00030
HS51 14 0 0 7 2 7 0.00001 0.00002 0.00004 0.00201 0.00025
CVXQP3_S 808 9 10 13 12 14 0.00052 0.00086 0.01384 0.00406 0.00455
CVXQP3_L 82481 - 10 - 18 24 - 21.84589 - 2.47151 106.68357
STADAT3 31997 16 54 28 26 22 0.01792 0.10851 0.07749 0.04597 0.07562
S268 40 5 10 15 12 15 0.00002 0.00005 0.00007 0.00239 0.00047
QSTANDAT 5030 24 20 11 18 18 0.00456 0.00757 0.03674 0.00550 0.05588
EXDATA 1137750 16 21 19 11 13 4.14068 5.91543 41.02600 0.93205 0.88089
QGROW7 3550 19 21 16 17 21 0.00292 0.00498 0.00441 0.00494 0.01086
DUAL4 3024 8 11 13 14 13 0.00105 0.00190 0.00228 0.00526 0.00334
QFORPLAN 5591 28 21 - 23 - 0.00554 0.00757 - 0.00635 -
QSCTAP3 12401 16 11 11 0 10 0.01539 0.01249 0.06928 0.00262 0.06155
QSHARE2B 828 19 15 24 18 21 0.00060 0.00086 0.00130 0.00313 0.00202
HS35MOD 12 9 11 11 9 13 0.00002 0.00004 0.00003 0.00203 0.00031
DUAL2 4796 8 10 13 13 8 0.00144 0.00237 0.00345 0.00495 0.00281
MOSARQP2 4775 8 9 27 16 13 0.00384 0.00567 0.02281 0.00931 0.01393
QPTEST 10 5 7 9 11 10 0.00002 0.00003 0.00003 0.00204 0.00030
AUG3D 9219 0 0 13 2 8 0.00148 0.00230 0.03595 0.00389 0.03126
QSIERRA 11557 22 26 38 18 30 0.01250 0.02299 0.04104 0.01296 0.05868
LISWET5 40002 8 9 - 13 24 0.01327 0.03012 - 0.05090 0.20424
LISWET3 40002 9 16 - 13 20 0.01723 0.05153 - 0.04864 0.15546
LISWET12 40002 - - - - 51 - - - - 0.61876
CVXQP3_M 8231 23 11 - 14 24 0.09018 0.05447 - 0.08596 0.55480
HS118 108 11 11 9 12 10 0.00008 0.00012 0.00010 0.00250 0.00051
QSHIP12S 28344 18 15 32 17 19 0.02393 0.02954 7.88479 0.01928 0.33531
DUALC1 1998 23 11 29 12 14 0.00088 0.00079 0.00184 0.00230 0.00084
CVXQP1_L 74982 - 10 - 13 34 - 20.36505 - 1.39803 127.34169
LISWET7 40002 11 - - 58 46 0.01622 - - 0.23969 0.53076
CONT-100 79596 11 9 22 12 20 0.27128 0.33981 0.67966 0.08552 0.35265
QSCFXM3 10369 31 32 35 29 43 0.01834 0.02726 0.06242 0.01993 0.10596
DPKLO1 1652 0 0 13 2 5 0.00016 0.00025 0.00173 0.00276 0.00111
KSIP 18871 14 14 21 17 11 0.00483 0.00848 0.00887 0.01248 0.00428
DUALC8 4076 14 9 27 10 17 0.00104 0.00143 0.00446 0.00241 0.00098
QSHIP08L 52050 14 15 28 19 17 0.09698 0.12749 11.94011 0.04549 0.47082
AUG3DQP 13092 11 11 31 18 17 0.01532 0.01932 0.08252 0.01974 0.10652
AUG2DQP 80000 11 13 - 19 - 0.07272 0.13726 - 0.05924 -
HS76 20 6 6 10 9 7 0.00002 0.00004 0.00004 0.00232 0.00054
QSCTAP2 9371 14 10 10 0 9 0.00771 0.00859 0.02937 0.00252 0.06080
BOYD1 745507 37 - - 30 - 0.93186 - - 0.31081 -
QCAPRI 3147 40 31 - 29 21 0.00544 0.00852 - 0.00698 0.01174
CONT-300 653093 12 10 - 12 - 13.65596 15.11321 - 0.84948 -
Table 9: Iterations and solver runtimes for SuiteSparse problems